1998 International Workshop on Vortex Dynamics in High-Temperature Superconductors Hachimantai, Iwate, Japan

June 21^{th}–June 26^{th}, 1998

Program

Main Sponsors Core Research for Evolutional Science and Technology (CREST) Program, Japan Science and Technology Coorporation (JST), Japan

Co-Sponsors Information, Science and Technology Division, Iwate Prefectural Government, Japan

Materials Science Division, Argonne National Laboratory, USA

Sunday, June 21^{th}

18^{00}  Arrival at Hachimantai Royal Hotel
19^{00}  Get Together Party and Dinner

Monday, June 22^{th}

S-Ia: Phase Diagram  Chair: Eli Zeldov, Weizmann Institute
8^{45}–9^{15}  Christophe Marcenat, Grenoble

Anisotropy of the specific heat of YBCO in the mixed state:

Scaling laws and vortex lattice melting in strong magnetic field (0 - 26 T)

9^{15}–9^{45}  Terukazu Nishizaki, Tohoku

Effects of the weak disorder on the vortex phase diagram up to 30 T in YBCO

9^{45}–10^{15}  Asle Sudbø, Trondheim

Vortex-loop disorder parameter in high-T_{c} superconductors

10^{15}–10^{30}  Discussion
10^{30}–11^{00}  Coffee break
S-Ib: Phase Diagram  Chair: Thomas Natterman, Köln
11^{00}–11^{30}  Dan Fuchs, Weizmann Institute

Possible new vortex matter phases in BSCCO

11^{30}–12^{00}  Daniel Lopez, Argonne

Longitudinal and transverse correlation with non-uniform current distributions

12^{00}–12^{30}  Hikaru Kawamura, KIT-Kyoto

Equilibrium simulation of longitudinal and transverse order of the three-

dimensional vortex glass

12^{30}–12^{45}  Discussion
12^{45}  Lunch
S-IIa: Phase Diagram  Chair: Mike Moore, Manchester
15^{00}–15^{30}  Xiao Hu, NRIM-Tsukuba

Structure and phase transition of the Josephson vortex system in high-T_{c} superconductors

15^{30}–16^{00}  Nai-Chang Yeh, Caltec

Effects of intrinsic and extrinsic pinning on the Josephson vortex phases and

dynamics of superconducting cuprates

16^{00}–16^{15}  Discussion
17^{15}–17^{45}  Coffee break
S-IIb: Phase Diagram  Chair: Lia Krusin-Elbaum, IBM - Yorktown Hights
16^{45}–18^{00}  Discussion
18^{30}  Dinner
Poster Preview  Chair: Wai Kwok, Argonne
20^{30}–22^{00}  1 minute × 80, with Beer

Tuesday, June 23^{th}

S-IIIa: Interplane Correlation Chair: Masashi Tachiki, NRIM-Tsukuba
8^{45}–9^{15}  Takasada Shibauchi, Hongo-Tokyo

Loss of interlayer phase coherence at the vortex lattice transition in layered SC

9^{15}–9^{45}  Alexei Koshelev, Argonne

Weak Josephson coupling as a probe of vortex state in layered Superconductors

9^{45}–10^{30}  Discussion
10^{30}–11^{00}  Coffee break
S-IIIb: Interplane CorrelationChair: Marcin Konczykowski, Palaiseau
11^{00}–11^{30}  Kees van der Beek, Palaiseau

Thermodynamics and vortex pinning by columnar defects in HTSC single crystals

11^{30}–12^{00}  Nikolai Morozov, Los Alamos

Evidences of vortex line recoupling in irradiated Bi-2212 from transport measurement

12^{00}–12^{30}  Ryoko Sugano, HARL-Hitachi

Monte Carlo study of vortex state in heavy ion irradiated Bi-2212

12^{30}–12^{45}  Discussion

12^{45}  Lunch

Poster Session I  Chair: Yuji Matsuda, ISSP-Tokyo
15^{00}–18^{00}  Poster Session I
18^{30}  Dinner
S-IV: Imaging of Vortics Chair: David Bishop, Bell Labs - Lucent Tech.
20^{30}–21^{00}  Akira Tonomura, HARL-Hitachi

Direct observation of the vortex motion at the depinning threshold in high-T_{c} Bi-2212 thin films by Lorenz microscope

21^{00}–21^{15}  Peter Kes, Leiden (short presentation)

Vortex motion probed by STM in NbSe_{2} in the presence of both point

and correlated disorders

21^{15}–21^{30}  Masanori Hangyo, Osaka (short presentation)

Ultrafast optical response and terahertz radiation

from superconducting YBCO films

21^{30}–22^{00}  Discussion

Wednesday, June 24^{th}

S-Va: Imaging of Vortices Chair: Mikhail Indenbom, ISSP-Chernogolovka
8^{45}–9^{15}  Vitalii Vlasko -Vlasov, Argonne

Peak effect in Bi and Tl cuprates - a disorder driven phase transition in the FLL

9^{15}–9^{45}  Michael Koblischka, SRL-ISTEC

Turbulant relaxation of magnetic fields in HTSC (tentative)

9^{45}–10^{30}  Discussion
10^{30}–11^{00}  Coffee break
S-Vb: Imaging of Vortices Chair: Tsuyoshi Tamegai, Hongo-Tokyo
11^{00}–11^{30}  Ted Forgan, Birmingham

Vortex structures in layered superconductors YBCO, Sr_{2}RuO_{4} and BSCCO

11^{30}–11^{45}  Jun-ichi Suzuki, JAERI-Tohkai (short presentation)

Small-angle neutron scattering study of Josephson vortices

in a superconductor Bi_{2}Sr_{2}CaCu_{2}O_{8+δ}

11^{45}–12^{15}  Ahmet Oral, Oxford

Direct observation of vortex lattice melting in BSCCO single crystals:

A scanning Hall probe microscopy study

12^{15}–12^{30}  Discussion

12^{45}  Lunch around Hachimantai area

Afternoon  Excursion + Dinner in Morioka City

21^{30}  Return to the Hotel

Thursday, June 25^{th}

S-VIa: Dynamics of Vortices Chair: Aharon Kapitulnik, Stanford
8^{45}–9^{15}  Flavio Pardo, Bell Lab., Lucent Tech.

Dynamical properties of the vortex lattice as revealedv by Bitter magnetic decoration

9^{15}–9^{45}  Ulrich Welp, Argonne

Driven vortex states and relaxation in single crystal Y124

9^{45}–10^{15}  Atsutaka Maeda, Komaba-Tokyo

Local vortex density fluctation in driven vortex systems in both the liquid and solid phases

10^{15}–10^{30}  Discussion
10^{30}–11^{00}  Coffee break
S-VIb: Dynamics of Vortices Chair: Hugo Safar, Illinois-Chicago
11^{00}–11^{30}  Franco Nori, Michigan

Non-equilibrium dynamic vortex phases

11^{30}–12^{00}  Thierry Giamarchi, Paris-Sud

Moving glass theory of driven lattices with disorder

12^{00}–12^{30}  Valerii Vinokur, Argonne

Glassy motion of vortex lattices

12^{30}–12^{45}  Discussion

12^{45}  Lunch

Poster Session II  Chair: Kazuo Hirata, NRIM-Tsukuba
15^{00}–17^{30}  Poster Session II
18^{00}  Dinner (Japanese Style)
S-VII: Discussion  Chair: George Crabtree, Argonne
20^{30}–22^{00}  Discussion on selected topics

Friday, June 26^{th}

S-VIIIa: Electronic State Chair: Koichi Kitazawa, Hongo-Tokyo
8^{45}–9^{15}  Øystein Fischer, Genevé

Scanning tunneling spectrospcopy of vortices in high temperature superconductors

9^{15}–9^{45}  Kazushige Machida, Okayama

Vortex core structure in s-wave and d-wave superconductors

9^{45}–10^{15}  Vadim Geshkenbein, ETH-Zürich

Temperature dependence of H_{c2} line in overdoped Tl_{2}Ba_{2}CuO_{6+δ}

10^{15}–10^{30}  Discussion
10^{30}–11^{00}  Coffee break
S-VIIIb: Pinning  Chair: John Clem, Iowa State
11^{00}–11^{30}  Yosi Yeshurun, Bar Ilan

Local voltage-current characteristics in high-T_{c} superconductors

11^{30}–12^{00}  Alexander Gurevich, Wisconsin-Madison

Mechanisms of current transport and vortex dynamics and pinning on grain boundaries

12^{00}–12^{30}  Gianni Blatter, ETH-Zürich

Thermal suppression of strong pinning

12^{30}–12^{45}  Discussion and Closing
12^{45}  Lunch
13^{30}  Departure